OpenStudy (anonymous):

HELP ME PLZ!! Beau can process 36 envelopes for a mailing in 24 minutes. At this rate, how many minutes will it take him to process 120 envelopes?

jimthompson5910 (jim_thompson5910):

36/24 = 120/x solve for x

OpenStudy (anonymous):

oh ok so 80

OpenStudy (anonymous):

thank you!!!!

jimthompson5910 (jim_thompson5910):

yep x = 80

OpenStudy (anonymous):

can you help me with another question @jim_thompson5910 please??

OpenStudy (anonymous):

STUDENTS AT A UNIVERSITY CAMPUS 22 or younger =44.7 23–25 years =25.2 26–35 years= 20.1 36 or older = 10.0 The table above shows the percent of students in different age groups at a university campus. The total number of students at the campus is 8,987. Approximately how many of the students at the campus are of age 25 years or younger?